# Synthetic Turbulence via an Instanton Gas Approximation

### v1

@@ IV.3 Including Gaussian fluctuations
[…]
@@ V Monte Carlo simulations of the instanton gas
Figure 6: Gramian $g$, as defined in Eq. (13), as a function of the gradient value $a=\partial_{x}u(0,t_{0})$, for the transformation onto the collective coordinates $(a,x_{0})$ of the instantons. Each open dot in the figure corresponds to one numerically computed instanton configuration as described in Sec. III. Note the double-log scale of the figure; in the shaded region around $a=0$, a linear horizontal scale is used instead. In particular, for large negative gradient values, we see that $g$ is almost constant.
For the case of interacting instantons, with joint PDF (11) for $2n_{\text{I}}$ random parameters (instanton positions and gradients), the density no longer factorizes. Hence one needs to use e.g. a MCMC based approach to sample from the target density (11), which is known up to a normalization constant. For intstance, one could use the classical Metropolis algorithm [100], where in each step of the chain, the proposal could consist of
@@
[…]
@@
×$\times\exp\left\{-\frac{1}{\sigma^{2}}\left(S\left(\sum_{j=1}^{n_{\text{I}}}u^{\text{I}}_{a_{j}^{\prime},x_{0j}^{\prime}}\right)-S\left(\sum_{j=1}^{n_{\text{I}}}u^{\text{I}}_{a_{j},x_{0j}}\right)\right)\right\}\,.$
@@
For this paper, for simplicity – as it requires less fine-tuning, and a robust and parallelized implementation is already available – we use the affine-invariant ensemble sampler of Ref. [101] as implemented by Ref. [102], instead of the Metropolis sampler, to sample the $6$-dimensional distribution (11) of the collective coordinates (for $n_{\text{I}}=3$). For each $\sigma^{2}$, we run $N_{\text{w}}=90$ interacting “walkers”, with independent initial states $\left\{a_{j}^{(k)},x_{0j}^{(k)}\right\}_{j=1}^{n_{\text{I}}}$, $k=1,\dots,N_{\text{w}}$, sampled from the non-interacting distribution (16), for $5000$ MCMC steps and measure the coordinate-wise autocorrelation time $\tau$ of the chain. A burn-in period of $2\tau$ is then used, and approximately independent samples (using all walkers) are subsequently obtained by subsampling the chains every $\tau/2$ steps. Since for the problem at hand, the velocity field (8) is $2\pi$-periodic in space and the PDF (11) hence only depends on $x_{0j}\text{ mod }2\pi$, we shift the instanton positions back to $[0,2\pi]$ every $100$ MCMC steps to avoid numerical errors arising from possibly large values of the $x_{0j}$’s.
@@ VI.2 Lognormal cascade model
To generate random fields with anomalous scaling but without the coherent structures, we consider the continuous wavelet cascade model introduced by Muzy [36]. The goal of this model, as a representative member of a class of synthetic turbulence models, is to mimic the multifractal statistical properties of turbulence, independently from the fluid’s equations of motion. The corresponding stochastic process $u$ from Ref. [36] is defined as the weak limit, when the small-scale cutoff $\ell$ tends to 00, of
@@
Here, $20<H<$ 2 is a hyperparameter that controls the smoothness of realizations (similar to the Hurst exponent for fractional Brownian motion), with smaller $H$ corresponding to rougher realizations. Randomness is introduced through the stochastic process $\omega_{s}(x)$, which is infinitely divisible in the scale $s$. The non-linear operation $\exp\{\omega_{s}(x)\}$ introduces intermittency and fulfills the multifractal scaling relation $\mathbb{E}[\exp\{p\,\omega_{s}(x)\}]=\left(L_{\mathrm{int}}/s\right)^{\phi(p)}$, where $\phi(p)$ denotes the cumulant generating function of $\omega_{s}(x)$. This construction constitutes the continuous counterpart to classical discrete multiplicative cascades [70]. The multifractal noise $\exp\{\omega_{s}(x)\}$ is translated scale-by-scale into the velocity signal $u(x)$ in Eq. (25) via convolution with the synthesizing wavelet function $\psi(x)$.
@@
Figure 7: Energy spectrum $E(k)=\mathbb{E}\lvert\hat{u}_{k}\rvert^{2}/2$ (top row) and one-point gradient PDF $\rho(a)$ with $a=\partial_{x}u(0,0)$ (bottom row) at different noise strengths $\sigma^{2}$ or Reynolds numbers for the random field ensembles listed in Sec. VII. Top row: As the noise strength increases, the characteristic $E(k)\propto k^{-2}$ spectrum of the Burgers Eq. (1) develops (dashed black lines for comparison). The DNS and phase-randomized DNS energy spectra exactly coincide by construction, and all other ensembles capture the energy spectrum’s $k$-dependence for the different $\sigma^{2}$’s relatively well. Bottom row: The DNS gradient PDF becomes more and more non-Gaussian, with a heavy left tail and light right tail, as $\sigma^{2}$ increases. This is mimicked correctly by the instanton gas without fluctuations, while the other ensembles yield quite different, symmetric distributions.
@@ VII Results
For each of these ensembles, we compute and analyze $N=10^{4}$ realizations of the random fields at all 15 noise strengths $\sigma^{2}\in[10^{-2},3\cdot 10^{5}]$ from Sec. III, or the corresponding equivalent ensembles. For a more detailed comparison of some statistical properties, we will use three representative noise strengths or (DNS) Reynolds numbers of $\sigma^{2}\in\{1.17\cdot 10^{-1},1.85\cdot 10^{2},2.51\cdot 10^{4}\}$ and $\text{Re}\in\{1.03,32.29,192.21\}$, respectively.
@@ VII.1 Eulerian statistics
This picture is qualitatively reproduced by the instanton gas ensembles (columns 2 and 3 of Fig. 1), as could be expected from Fig. 4. We see that, as remarked in Sec. IV.2, the non-interacting ensemble always generates fields with (here) three instantons by construction. Depending on the realization of the independently and uniformly positions $x_{0j}$ of the instantons, this may give “too many” shocks at high Reynolds numbers, compared to the DNS fields. In contrast to this, the interacting ensemble, through volume exclusion effects of the interactions $\Delta S$ in Eq. (14), can lead to only effectively one (Fig. 1, $\sigma^{2}=25118.864$, red curve) or two (Fig. 1, $\sigma^{2}=25118.864$, orange and blue curves) instantons being present, which appears in principle closer to DNS field realizations. The appearance of effective single-instanton configurations and regularly spaced two-instanton configurations in the interacting ensemble suggests that the effect of interactions is relatively strong in our particular setup. Going to larger domain sizes, or equivalently shorter forcing correlation lengths, could lead to a more dilute gas with a larger number of weakly interacting instantons.
@@
Figure 8: Structure function exponents $\zeta_{p}$, determined from log-log linear fits to $S_{p}(r)\propto r^{\zeta_{p}}$, as defined in Eq. (28), in the inertial range, for the random field ensembles at different noise strengths $\sigma^{2}$ or Reynolds numbers. The linear fits were performed in the range $r\in[0.1,1]$ for all ensembles and noise variances $\sigma^{2}$ for simplicity. In ensembles 4. and 6., we show only exponents until $p=4$, as higher-order structure functions did not display any clear scaling behavior in the fit range. The structure function exponents determined from DNS approach the shock scaling $\zeta_{p}=\min\{1,p\}$ as $\sigma^{2}$ increases, and this is reproduced by the instanton gas ensembles 2. and 3.
@@ VII.2.1 In one dimension
$\begin{cases}\frac{\mathrm{d}x}{\mathrm{d}t}=v(t)\,,\quad&x(0)=0\,,\\ \frac{\mathrm{d}v}{\mathrm{d}t}=u(x(t))\,,\quad&v(0)=v_{0}\,,\end{cases}$ (29)
Here, we start the particles at the left end $x=0$ of the box $[0,L_{\text{box}}]$, and vary their initial velocity $v_{0}\in[0.2,5]$. Note that this one-dimensional situation is necessarily quite different from the magnetic Lorentz force in three dimensions, in that the force in Eq. (29) can change the magnitude of the particles’ velocities, and is hence more comparable to an electric field. Nevertheless, particles can still be “captured” (similar to magnetic mirrors in MHD) or accelerated by shock-like structures. The setup is illustrated as a phase portrait in Fig. 9. We see that the quasi-shocks act as centers in the phase portrait. If the particle has a small initial velocity $v_{0}$ at $x=0$, it is captured by the first shock. If instead it starts above the separatrix, it will traverse the domain, but the time it takes will still be affected by the structure of the random field (less so as $v_{0}$ increases).
@@
Figure 10: Comparison of normalized average time for particles to cross the domain $r_{t}=\mathbb{E}\left[t_{\text{arrival}}\mid t_{\text{arrival}}<\infty\right]/t_{\text{ballistic}}$ (dots), where $t_{\text{ballistic}}=L_{\text{box}}/v_{0}$, and fraction of particles that manage to cross the domain $r_{N}=N_{\text{arrival}}/N$ (crosses) from $x=0$ to $x=L_{\text{box}}$, for the different random field ensembles when varying the noise strength $\sigma^{2}$ or Reynolds number. All $N=10^{4}$ fields per ensemble are normalized to ensemble unit root-mean square velocity for this experiment, and $r_{t},r_{N}$ are shown as a function of the initial velocity $v_{0}$, cf. Fig. 9. While all ensembles yield similar results at small $\sigma^{2}$, the effect of increasing the Reynolds number is different between the ensembles, notably between those that include coherent structures and those that do not. For a better comparison, the insets show the quotient of the respective ensemble’s $r_{t}$ and $r_{N}$, relative to the DNS result.
@@ VIII Conclusion and outlook
Among the different instanton ensembles considered here, already a canonical ensemble of instantons without interactions or fluctuations seems be successful in capturing the DNS statistics quite well. This will not necessarily be the case in other systems beyond the one-dimensional Burgers equation. We have seen that including interactions among instantons, which can incorporate volume exclusion effects and effectively remove some of the instantons, only yields small modifications of the Eulerian and Lagrangian statistics we analyzed, and in some cases even results in worse agreement with DNS than in the non-interacting case. Similarly, we have discussed the inclusion of independent Gaussian fluctuations around individual instantons. While Gaussian fluctuations considerably improve the theoretical one-point gradient PDF prediction as shown in Fig. 5 even at high noise strengths $\sigma^{2}$, they tend to overpower the instantons in the random fields for instanton observable values $a_{j}$ that are in the core of the gradient PDF, such that the resulting fields look almost Gaussian and hence less and less like DNS fields at high $\sigma^{2}$.
@@
We show a preliminary proof of concept in Fig. 13. Here, instantons were calculated for the two-dimensional incompressible MHD equations with the current density $j=(\nabla\times\bm{B})_{z}=\partial_{x}B_{y}-\partial_{y}B_{x}$ at one point in the domain as observable, using an optimization procedure similar to that described in this paper and following Ref. [59]. As a placeholder for the sampling procedures described in this paper, we used the one-point current density PDF obtained from a pseudo-spectral DNS run of the stochastically forced 2D MHD equations (see Fig. 13 (left) for a DNS field snapshot) to independently draw the different strengths of the current density of the instantons  [^3] Note that this is not consistent with Eq. (16) for non-interacting instantons, which instead includes the instanton action and Gramian for sampling the observable values. For the purpose of this qualitative outlook, we content ourselves with drawing from the one-point DNS PDF without Gramian, since it requires computing fewer instantons to generate a single snapshot – we draw observable values from the DNS PDF, and compute instantons just for those values.. Each instanton is randomly rotated and shifted, before being summed to form the final turbulent current density field in Fig. 13 (center). In addition, Fig. 13 (right) shows a turbulent field resulting from subsequent passive advection of the instanton gas, motivated by the findings discussed in Ref. [115]. To achieve this, a large-scale Gaussian random velocity field with a few Fourier modes ($\lVert k\rVert<6$) was used to advect the stream function corresponding to the instanton gas realization from Fig. 13 (center). Many aspects still require further investigation, e.g. whether there is a need for self-avoiding superposition of current sheets, whether there are further constraints besides numerical issues to produce fast methods applicable on large-scale domains, and whether the effect of passive advection and entanglement of structures can be achieved in a more self-consistent way, e.g. through interactions and fluctuations. Nevertheless, on a qualitative level, the field realization in Fig. 13 (right) already displays similar structures as the DNS snapshot.
@@ Appendix A Estimation of hyperparameters for the log-normal continuous cascade model
[…]

### v2

@@ IV.3 Including Gaussian fluctuations
with random normal coefficients $Z_{jk}\overset{\text{iid}}{\sim}{\cal N}(0,1)$ within a Karhunen–Loève expansion [101] of the conditioned Gaussian fluctuations at time $t_{0}$. Here, $\nu_{a}^{(k)}\geq 0$ and $\delta v_{a,x_{0}}^{(k)}$ are the eigenvalues and orthonormal eigenfunctions, respectively, of the integral operator $K_{a,x_{0}}$ acting as $(K_{a,x_{0}}v)(x)=\int_{0}^{2\pi}{\cal C}_{a,x_{0}}(x,x^{\prime})v(x^{\prime})\mathrm{d}x^{\prime}$, which is defined through the final-time covariance function
@@ V Monte Carlo simulations of the instanton gas
Figure 7: Gramian $g$, defined in Eq. (14), as a function of the gradient value $a=\partial_{x}u(0,t_{0})$, for the transformation onto the collective coordinates $(a,x_{0})$ of the instantons. Each open dot corresponds to one numerically computed instanton configuration. Note the double-log scale of the figure; in the shaded region around $a=0$, a linear horizontal scale is used instead. For large negative gradient values, $g$ is almost constant.
For the case of interacting instantons, with joint PDF (12) for $2n_{\text{I}}$ random parameters (instanton positions and gradients), the density no longer factorizes. Hence one needs to use e.g. a MCMC based approach to sample from the target density (12), which is known up to a normalization constant. For instance, one could use the classical Metropolis algorithm [102], where in each step of the chain, the proposal could consist of
@@
    changing the position $x_{0\hat{j}}\to x_{0\hat{j}}^{\prime}=x_{0\hat{j}}+\delta_{x_{0}}\cdot Z_{1}$ and gradient strength $a_{\hat{j}}\to a_{\hat{j}}^{\prime}=a_{\hat{j}}+\delta_{a}\cdot Z_{2}$ with $Z_{1},Z_{2}\overset{\text{iid}}{\sim}{\cal N}(0,1)$,
@@
$\times\exp\left\{-\frac{1}{\sigma^{2}}\left(S\left(\sum_{j=1}^{n_{\text{I}}}u^{\text{I}}_{a_{j}^{\prime},x_{0j}^{\prime}}\right)-S\left(\sum_{j=1}^{n_{\text{I}}}u^{\text{I}}_{a_{j},x_{0j}}\right)\right)\right\}\,.$
@@
For this paper, for simplicity – as it requires less fine-tuning, and a robust and parallelized implementation is already available – we use the affine-invariant ensemble sampler of Ref. [103] as implemented by Ref. [104], instead of the Metropolis sampler, to sample the $6$-dimensional distribution (12) of the collective coordinates (for $n_{\text{I}}=3$). For each $\sigma^{2}$, we run $N_{\text{w}}=90$ interacting “walkers”, with independent initial states $\left\{a_{j}^{(k)},x_{0j}^{(k)}\right\}_{j=1}^{n_{\text{I}}}$, $k=1,\dots,N_{\text{w}}$, sampled from the non-interacting distribution (17), for $5000$ MCMC steps and measure the coordinate-wise autocorrelation time $\tau$ of the chain. A burn-in period of $2\tau$ is then used, and approximately independent samples (using all walkers) are subsequently obtained by subsampling the chains every $\tau/2$ steps. Since for the problem at hand, the velocity field (9) is $2\pi$-periodic in space and the PDF (12) hence only depends on $x_{0j}\text{ mod }2\pi$, we shift the instanton positions back to $0,2\pi]$ every $100$ MCMC steps to avoid numerical errors arising from possibly large values of the $x_{0j}$’s.
@@ VI.2 Lognormal cascade model
To generate random fields with anomalous scaling but without the coherent structures, we consider the continuous wavelet cascade model introduced by Muzy [36]. The goal of this model, as a representative member of a class of synthetic turbulence models, is to mimic the multifractal statistical properties of turbulence, independently from the fluid’s equations of motion. The corresponding stochastic process $u$ from Ref. [36] is defined as the weak limit, when the small-scale cutoff $\ell$ tends to 0, of
@@
Here, $0<H<2$ is a hyperparameter that controls the smoothness of realizations (similar to the Hurst exponent for fractional Brownian motion), with smaller $H$ corresponding to rougher realizations. Randomness is introduced through the stochastic process $\omega_{s}(x)$, which is infinitely divisible in the scale $s$. The non-linear operation $\exp\{\omega_{s}(x)\}$ introduces intermittency and fulfills the multifractal scaling relation $\mathbb{E}[\exp\{p\,\omega_{s}(x)\}]=\left(L_{\mathrm{int}}/s\right)^{\phi(p)}$, where $\phi(p)$ denotes the cumulant generating function of $\omega_{s}(x)$. This construction constitutes the continuous counterpart to classical discrete multiplicative cascades [70]. The multifractal noise $\exp\{\omega_{s}(x)\}$ is translated scale-by-scale into the velocity signal $u(x)$ in Eq. (26) via convolution with the synthesizing wavelet function $\psi(x)$.
@@
Figure 8: Energy spectrum $E(k)=\mathbb{E}\lvert\hat{u}_{k}\rvert^{2}/2$ (top row) and one-point gradient PDF $\rho(a)$ with $a=\partial_{x}u(0,0)$ (bottom row) at different noise strengths $\sigma^{2}$ or Reynolds numbers for the random field ensembles listed in Sec. VII. The dashed black lines in the top row indicate $E(k)\propto k^{-2}$.
@@ VII Results
For each of the ensembles listed above, we compute and analyze $N=10^{4}$ realizations of the random fields at all 15 noise strengths $\sigma^{2}\in[10^{-2},3\cdot 10^{5}]$ from Sec. III, or the corresponding equivalent ensembles. For a more detailed comparison of some statistical properties, we will use three representative noise strengths or (DNS) Reynolds numbers of $\sigma^{2}\in\{1.17\cdot 10^{-1},1.85\cdot 10^{2},2.51\cdot 10^{4}\}$ and $\text{Re}\in\{1.03,32.29,192.21\}$, respectively.
@@ VII.1 Eulerian statistics
This picture is qualitatively reproduced by the instanton gas ensembles (columns 2 and 3 of Fig. 1), as could be expected from Fig. 4. We see that, as remarked in Sec. IV.2, the non-interacting ensemble always generates fields with (here) three instantons by construction. Depending on the realization of the independently and uniformly positions $x_{0j}$ of the instantons, this may give “too many” shocks at high Reynolds numbers, compared to the DNS fields. In contrast to this, the interacting ensemble, through volume exclusion effects of the interactions $\Delta S$ in Eq. (15), can lead to only effectively one (Fig. 1, $\sigma^{2}=25118.864$, red curve) or two (Fig. 1, $\sigma^{2}=25118.864$, orange and blue curves) instantons being present, which appears in principle closer to DNS field realizations. The appearance of effective single-instanton configurations and regularly spaced two-instanton configurations in the interacting ensemble suggests that the effect of interactions is relatively strong in our particular setup. Going to larger domain sizes, or equivalently shorter forcing correlation lengths, would lead to a more dilute gas with a larger number of weakly interacting instantons as discussed in Sec. IV.2.
@@
Figure 9: Structure function exponents $\zeta_{p}$, determined from log-log linear fits to $S_{p}(r)\propto r^{\zeta_{p}}$, as defined in Eq. (29), in the inertial range, for the random field ensembles at different noise strengths $\sigma^{2}$ or Reynolds numbers. The linear fits were performed in the range $r\in[0.1,1]$ for all ensembles and noise variances $\sigma^{2}$. In ensembles 4. and 6., we show only exponents until $p=4$, as higher-order structure functions did not display any clear scaling behavior in the fit range.
@@ VII.2.1 In one dimension
[…]
Here, we start the particles at the left end $x=0$ of the box $0,L_{\text{box}}]$, and vary their initial velocity $v_{0}\in[0.2,5]$. Note that this one-dimensional situation is necessarily quite different from the magnetic Lorentz force in three dimensions, in that the force in Eq. (30) can change the magnitude of the particles’ velocities, and is hence more comparable to an electric field. Nevertheless, particles can still be “captured” (similar to magnetic mirrors in MHD) or accelerated by shock-like structures. The setup is illustrated as a phase portrait in Fig. 10. We see that the quasi-shocks act as centers in the phase portrait. If the particle has a small initial velocity $v_{0}$ at $x=0$, it is captured by the first shock. If instead it starts above the separatrix, it will traverse the domain, but the time it takes will still be affected by the structure of the random field (less so as $v_{0}$ increases).
@@
Figure 11: Comparison of normalized average time for particles to cross the domain $r_{t}=\mathbb{E}\left[t_{\text{arrival}}\mid t_{\text{arrival}}<\infty\right]/t_{\text{ballistic}}$ (dots), where $t_{\text{ballistic}}=L_{\text{box}}/v_{0}$, and fraction of particles that manage to cross the domain $r_{N}=N_{\text{arrival}}/N$ (crosses) from $x=0$ to $x=L_{\text{box}}$, for the different random field ensembles when varying the noise strength $\sigma^{2}$ or Reynolds number. All $N=10^{4}$ fields per ensemble are normalized to ensemble unit root-mean square velocity for this experiment, and $r_{t},r_{N}$ are shown as a function of the initial velocity $v_{0}$, cf. Fig. 10. The insets show the quotient of the respective ensemble’s $r_{t}$ and $r_{N}$, relative to the DNS result.
@@ VIII Conclusion
Among the different instanton ensembles considered here, already a canonical ensemble of instantons without interactions or fluctuations seems be successful in capturing the DNS statistics quite well. This will not necessarily be the case in other systems beyond the one-dimensional Burgers equation. We have seen that including interactions among instantons, which can incorporate volume exclusion effects and effectively remove some of the instantons, only yields small modifications of the Eulerian and Lagrangian statistics we analyzed, and in some cases even results in worse agreement with DNS than in the non-interacting case. Similarly, we have discussed the inclusion of independent Gaussian fluctuations around individual instantons. While Gaussian fluctuations considerably improve the theoretical one-point gradient PDF prediction as shown in Figs. 5 and 6 even at high noise strengths $\sigma^{2}$, they tend to overpower the instantons in the random fields for instanton observable values $a_{j}$ that are in the core of the gradient PDF, such that the resulting fields look almost Gaussian and hence less and less like DNS fields at high $\sigma^{2}$.
@@ IX Outlook: 2D MHD turbulence
We show a preliminary proof of concept in Fig. 14. Here, instantons were calculated for the two-dimensional incompressible MHD equations with the current density $j=(\nabla\times\bm{B})_{z}=\partial_{x}B_{y}-\partial_{y}B_{x}$ at one point in the domain as observable, using an optimization procedure similar to that described in this paper and following Ref. [59]. As a placeholder for the sampling procedures described in this paper, we used the one-point current density PDF obtained from a pseudo-spectral DNS run of the stochastically forced 2D MHD equations (see Fig. 14 (left) for a DNS field snapshot) to independently draw the different strengths of the current density of the instantons  [^4] Note that this is not consistent with Eq. (17) for non-interacting instantons, which instead includes the instanton action and Gramian for sampling the observable values. For the purpose of this qualitative outlook, we content ourselves with drawing from the one-point DNS PDF without Gramian, since it requires computing fewer instantons to generate a single snapshot – we draw observable values from the DNS PDF, and compute instantons just for those values.. At this point, we would like to note that in general, spontaneous breaking of continuous symmetries of the instanton solutions can occur, cf. [59]. This is also the case for the present example of 2D MHD turbulence, where rotational symmetry is broken by the instanton solutions for the current density. The parameters that label the submanifold of possible instanton solutions (here, a single angle) then have to be included as additional collective coordinates for the instanton gas.
@@
Therefore, each instanton is randomly rotated and shifted, before being summed to form the final turbulent current density field in Fig. 14 (center). In addition, Fig. 14 (right) shows a turbulent field resulting from subsequent passive advection of the instanton gas, motivated by the findings discussed in Ref. [117]. To achieve this, a large-scale Gaussian random velocity field with a few Fourier modes ($\lVert k\rVert<6$) was used to advect the stream function corresponding to the instanton gas realization from Fig. 14 (center). Many aspects still require further investigation, e.g. whether there is a need for self-avoiding superposition of current sheets, whether there are further constraints besides numerical issues to produce fast methods applicable on large-scale domains, and whether the effect of passive advection and entanglement of structures can be achieved in a more self-consistent way, e.g. through interactions and fluctuations. Nevertheless, on a qualitative level, the field realization in Fig. 14 (right) already displays similar structures as the DNS snapshot.
@@ Appendix A Estimation of hyperparameters for the log-normal continuous cascade model
$\sigma^{2}$ $\ell$ $H_{\mathrm{actual}}$ $H_{\mathrm{input}}$ $\mu_{\mathrm{actual}}$ $\mu_{\mathrm{input}}$